# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v2

@@ 4.4 Task Definitions
Figure 3: TIPO’s task flow within a single generation cycle. Starting from any node, each arrow represents a sequential extension step, with other prompt used as metadata if provided. This task design enables efficient and flexible prompt extension across multiple tasks.
TIPO defines three core tasks: extending tag sequences, extending NL prompts, and generating refined NL prompts. Each task is represented by an arrow in Figure 3, illustrating how TIPO can use tag and NL prompts interchangeably as input sources to generate detailed outputs within a single generation cycle. When an arrow cycle is established, TIPO can complete multiple extension within a single generation pass.
For example, TIPO can start from a short NL prompt to produce a detailed tag sequence, or it can combine tasks like short_to_tag and tag_to_long to create a composite task, short_to_tag_to_long, which generates an extended NL prompt from an initial short NL input via tag generation. Further details on specific task sequences are provided in Appendix.
4.5 An Example Scenario and Workflow
Figure 4: An example scenario of TIPO workflow. (a) A generated image and prompts. (b) Prompt optimization of TIPO, starting from simple user input $p_{s}$ to detailed final output $p_{d}$. The shading from gray to light sky blue represents an increase in context richness in the prompt.
Figure 4 illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available. In such cases, TIPO processes each input type separately to maintain clarity and coherence. During each generation cycle, one input type is treated as the primary prompt which is target prompt for extending or refining, while the other types are used as <Content> of metadata to enrich the context. Figure 4 provides an example of this setup, where:
@@ 4.5 An Example Scenario and Workflow
$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\verb|<meta>|=\emptyset{}\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, ...}\right\}\end{array}$
@@
    short_to_tag: TIPO uses $T_{s}$ as the primary prompt to generate a detailed tag sequence $T_{d}$
@@
    short_to_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed prompt.
@@
    TIPO aggregates $T_{d}$, $S_{d}$, and any additional metadata to construct a context rich prompt, $p_{d}$.
This iterative process allows TIPO to build prompts that are both detailed and contextually aligned with user inputs, through short_to_tag, tag_to_long, and short_to_tag_to_long tasks within a single generation cycle.
4.6 Model Architecture
TIPO utilizes a causal autoregressive language model framework as its prompt extension function, $\mathcal{E}$, to iteratively extend and refine user-provided prompts into more detailed, content-rich prompts $p_{d}$ based on simpler inputs $p_{s}$. Our implementation leverages the LLaMA architecture [61, 62, 3], employing models with 200M and 500M parameters.
@@ 4.6 Model Architecture
Throughout training, non-padding tokens were used for token count calculations, with the effective count reflecting the relatively short and variable length of data entries.
4.7 Training Process
In each training pass, TIPO applies dataset augmentation by selecting tasks from all possible task combinations such as short_to_tag_to_long, including extending tag sequences, NL prompt extension, and refined NL prompt generation. This random selection across all task combinations significantly broadens the dataset’s effective scope, enhancing model generalization without additional manual engineering or runtime tuning.
(a) Direct LLM
(b) Prompt DB
(c) RL (Promptist)
(d) TIPO (Ours)
Figure 5: In-domain test regard to scenery test examples from 4 different methods.
@@ 4.7 Training Process
Figure 6: Comparison of generated images: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated long prompt, and (d) TIPO-enhanced truncated prompt. TIPO adds detail and maintains variety, yielding coherent images even from minimal prompts.
5 Evaluation
5.1 Experimental Setup
In this section, we detail the experimental setup used to evaluate prompt optimization framework. To assess the quality of generated prompts and resulting images, we employed the following evaluation metrics.
FDD (Frechet Dino Distance).
Traditionally, the Frechet Inception Distance (FID) [28] has been used to measure the distribution distance between datasets and generated images. However, recent works have shown that FID is not always aligned with human preferences [58]. Therefore, we measure the Frechet Distance using DinoV2 [46] outputs. In our experiments, we used the pretrained ViT-g [20] model from DinoV2 and used the normalized embedding vector to measure the Frechet Distance result.
Aesthetic Score.
We compute the Aesthetic Score using the Aesthetic Predictor V2.5 [19]. Higher aesthetic scores indicate images that are perceived as more visually appealing, well-composed, and of higher quality. This metric aims to quantify the subjective visual appeal of generated images, with higher scores suggesting outputs that are more likely to be judged favorably by human viewers.
AI Corrupt Score.
The AI Corrupt Score is obtained from the AICorruptMetrics [43] in sdeval [44]. This metric is trained on AI-generated images with human-annotated “corrupt or not” labels. A higher value indicates a higher likelihood of the image being “correct” or “complete”.
Vendi Score.
To quantify the diversity of generated images, we employ the Vendi Score metric based on DinoV2 [46] embeddings. The computation involves calculating a similarity matrix $K$ using cosine similarity between image embeddings, followed by computing the von Neumann entropy of the normalized matrix $K/n$, adhering to the standard Vendi Score methodology. This metric provides a quantitative measure of the distributional diversity within a set of generated images. Higher Vendi Scores indicate greater inter-image diversity, reflecting the model’s capability to generate varied and distinct outputs.
Extension Functions.
We used both GPT-4o-mini [2] and LLaMA-3.2-1B [3] as Direct LLM models for prompt generation, balancing performance and efficiency. GPT-4o-mini, as a lightweight closed-source model, offers speed and affordability, while LLaMA-3.2-1B represents a similar scale to TIPO in open-source models. Also, Prompt DB [17] and Promptist [26] serve as baseline methods for prompt extension.
5.2 In-domain test regard to scenery tag
To evaluate the model’s capability to generate diverse and accurate outputs, we chose the scenery tag, which includes both natural landscapes and human-made structures. In this test, we randomly select 32,768 entries from Danbooru2023 [65, 45] that include the “scenery” tag. We set up the following prompt:
<meta> + scenery + $\mathcal{E}$>
Then, we apply extension functions to extend only the scenery, $T_{s}$, as the <Content> tag. We generate one image from each prompt using the Kohaku-XL-zeta [^2] https://huggingface.co/KBlueLeaf/Kohaku-XL-Zeta model, which is a fine-tuned SDXL [48] model on the Danbooru dataset [45].
Metric Original LLM Prompt DB Promptist TIPO (Ours)
FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
Table 1: Metric results on in-domain test. Bold values indicate the best performance, and underlined values indicate the second-best. Original refers to images generated using the original text captions; LLM uses GPT-4o-mini [2]; the remaining columns represent results from Prompt DB [25], Promptist [26], and TIPO, respectively.
As shown in Table 1, TIPO-generated prompts improve FDD scores by aligning images more closely with the original dataset distribution. Enhanced AI Corrupt scores show that TIPO reduces image corruption, creating more coherent outputs by replacing vague descriptions (e.g., detailed background) with specific elements (e.g., garden filled with forget-me-nots and blue butterflies). The direct LLM approach highlights a key issue in prompt engineering: it achieves high aesthetic scores but compromises FDD, reducing the diversity of images.
We provide more comparative visual result in Figure 5. The results reveal significant biases in Prompt DB  [25] and RL-based [26] approaches. PromptDB and Direct-LLM method exhibits notable bias, by generating mountain/lake scene majorly. Meanwhile, Promptist produces diverse near-future fantasy sceneries, however often produces corrupted images due to unreasonable vocabulary usage (4K, masterpiece, instagram), supposedly introduced by user prompts.
By contrast, TIPO consistently reproduces a balanced output distribution, incorporating both natural and architectural elements, effectively mitigating the biases observed in other methods.
(a) Original prompt best aesthetic
(b) GPT-4o-mini best aesthetic
(c) TIPO best aesthetic
Figure 7: Top-10% aesthetic score images generated by the SD-3.5-Large model [19]. Rows show samples from Original, Direct LLM-Modified prompts, and TIPO-extended prompts. Direct LLM images focus primarily on natural scenery, while generated image from TIPO-extended prompt display a wider variety, including urban scenes and human figures.
5.3 In-domain prompt generation test
We aim to test each models’ performance on regard to reconstructing / upsampling the given prompt, in given dataset distribution. By comparing original ground truth and generated outputs, we can directly compare FDD, aesthetic score, AI corrupt scores.
We randomly selected 10k entries each from [12] and [29]. For each entry, we extracted the short caption and the truncated long caption (limited to a maximum of two sentences). These captions serve as user input prompts $p_{s}$ for in-domain prompt generation test. Using these prompts, we generated single image per prompt with the SDXL-1.0-base [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 model [48].
Short Original LLM Prompt DB Promptist TIPO(Ours)
FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
Truncated Long Original LLM Prompt DB Promptist TIPO(Ours)
FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
Table 2: Performance comparison in the in-domain prompt generation test. Bold values indicate the best performance, and underlined values indicate the second-best. Direct LLM achieves the highest Aesthetic Score but struggles with OOD issues. TIPO provides a balanced outcome, excelling in AI Corrupt Score and maintaining strong alignment, making it ideal for practical applications.
Table 2 presents the results of our Short/Truncated caption test for prompt generation. Compared to in-domain test, Section 5.2, which might be considered slightly out of domain by SDXL model, we observed that prompts covered by this test produces meaningful, coherent images with sufficient detail level. FDD scores were degraded compared to original-ground truth prompts, not truncated long caption-which implies prompt extension functions resulted lack of diversity, or unmatched components in generation.
We observe that Direct LLM approach achieves the highest aesthetic scores by wide vocabulary with user preference matching, however results low FDD, AI corruption score, which implies unmatched vocabularies are being used in-prompt.
Original LLM Prompt DB Promptist TIPO(Ours)
Top 10% $\uparrow$ 10.98 7.84 12.17 10.12 17.33
Worst 10% $\uparrow$ 12.10 9.76 15.23 14.14 13.92
Table 3: Vendi Scores [23] for the top and worst 10% aesthetic images across prompt generation methods. Higher value indicate higher diversity of samples. TIPO achieves high score in the top 10% (17.33), showing greater visual diversity as well as worst 10%. Result suggests TIPO’s balanced and robust variety in generated images. Meanwhile, LLM has the lowest scores in both categories, indicating limited diversity and repetition.
5.4 Out of focus evaluation.
In this section, we analyze TIPO’s behavior in out-of-focus model, SD-3.5-Large [^4] https://huggingface.co/stabilityai/stable-diffusion-3.5-large built upon MMDiT architecture [21]. We compose original tag, natural language prompt created by GPT-4o-mini, tag+natural language extended prompt by TIPO (ours), then feed to SD-3.5-Large inference pipeline.
Analysis of Vendi Scores [23] (Table 3) reveals contrasting patterns: Direct LLM-generated images show limited diversity despite high aesthetic scores, while TIPO maintains conceptual diversity across all quality levels. However, TIPO’s training on diverse datasets [45, 65, 12, 29] introduces domain-specific elements that may not align with SD3.5L’s training distribution. This mismatch demonstrates our core finding: prompt generation models perform optimally when trained on domain-matched prompt corpora.
The balanced profile of TIPO suggests that TIPO can work as suitable solution for prompt upsampling task, which can accurately project trained prompt with few initial word input.
@@ 6 Conclusion
In this paper, we introduced TIPO (Text to Image with text Presampling for Prompt Optimization), a novel framework designed to enhance the quality of Text-to-Image (T2I) models through automatic prompt engineering. TIPO represents a significant advancement in the field of T2I generation, offering a versatile and efficient solution for prompt optimization. By automating the refinement and extension of user prompts. We also highlights the importance of aligning user inputs with underlying data distributions through the comparisons between different methods for prompt engineering. As generative models continue to evolve, frameworks like TIPO will be instrumental in unlocking their full potential, ensuring they meet diverse user needs with precision and creativity while pushing the boundaries of what’s possible in generative models.

### v3

@@ 4.3 Text Pre-sampling
Figure 3: TIPO’s task flow within a single generation. Starting from any node, each arrow represents a sequential extension step, with other prompts used as metadata if provided. This task design enables efficient and flexible prompt extension across multiple tasks.
We propose the core technique of TIPO, a flexible pre-sampling mechanism that divides the prompt optimization process into three interconnected subtasks: enriching tag sequences, extending NL prompts, and refining NL prompts. As shown in Figure 3, TIPO uses both tag and NL prompts interchangeably as input sources to generate detailed outputs in various forms. For instance, TIPO can begin with a short NL prompt to produce a detailed tag sequence, denoted as short_to_tag. Training data for the language model is constructed by concatenating different types of $p_{o}$ with task-specific metadata embedded into the text.
The individual tasks are defined as follows:
  • tag_to_long: Use tags as metadata to generate a new NL prompt.
  • long_to_tag: Use an NL prompt as metadata to extend a tag sequence.
  • short_to_tag: Use the simple prompt $p_{s}$ as metadata to extend a tag sequence.
  • short_to_long: Use a user-provided NL prompt as metadata to generate a refined, detailed prompt.
We also consider composite tasks that can be performed within a single forward pass. These tasks offer the language model a holistic view of the objective while reducing computational overhead:
  • short_to_tag_to_long: Use a user-provided NL prompt or tag sequence as metadata to generate a refined, detailed prompt.
  • short_to_long_to_tag: Use a user-provided NL prompt or generated NL prompt as metadata to extend a tag sequence.
  • tag_to_short_to_long: Use user-provided tags or NL prompts as metadata to generate a refined NL prompt.
We randomly select from the aforementioned tasks during training to enhance model generalization. By extensively training on these tasks, TIPO can seamlessly adapt to various input types, flexibly refining user input whether it consists of tags, short sentences, or long sentences. Figure 4 illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available. In such cases, TIPO processes each input type separately to maintain clarity and coherence:
@@
$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\texttt{metadata}=\emptyset\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, \ldots}\right\}\end{array}$
@@
    short_to_tag: TIPO uses $T_{s}$ as the primary prompt to generate a detailed tag sequence $T_{d}$.
@@
    short_to_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed output.
@@
    TIPO aggregates $T_{d}$, $S_{d}$, and any additional metadata to construct a context-rich prompt $p_{d}$.
This progressive process enables TIPO to build prompts that are both detailed and contextually aligned with the user’s input.
Figure 4: An example scenario of TIPO workflow. (a) A generated image and prompts. (b) Prompt optimization of TIPO, from simple user input $p_{s}$ to detailed final output $p_{d}$. The shading from gray to light sky blue represents an increase in context richness in the prompt.
4.4 Implementation Details
TIPO can utilize various types of causal autoregressive language models by design. In practice, we leverage the LLaMA architecture [64, 65, 3] for simplicity, with 200M and 500M parameters.
@@ 4.4 Implementation Details
5 Experiments
5.1 Experimental Settings
Baselines. We compare with state-of-the-art prompt optimization methods as follows:
  • GPT [2] series are powerful large language models capable of zero-shot prompt optimization. We use GPT-4o-mini for lower cost and better efficiency.
  • MagicPrompt [17] trains GPT-2 on high-quality prompts collected from stable-diffusion model users.
  • Promptist [26] optimizes user input into model-preferred prompts via reinforcement learning.
Evaluation Metrics. We employ four latest metrics FDD [61], Aesthetic Score [19], AI Corrupt Score [45], and Vendi Score [22] to measure the quality of generated images. Specifically, FDD (Frechet Dino Distance) measures the image fidelity using the similarity between generated and dataset images via DinoV2 features [47], which better aligns with human perception than traditional FID [28]. Aesthetic Score is computed via Aesthetic Predictor V2.5 [19], quantifying visual appeal, composition quality, and artistic merit. AI Corrupt Score detects technical flaws in generated images by identifying visual artifacts. Vendi Score quantifies image diversity by calculating the von Neumann entropy from a normalized cosine similarity matrix using DinoV2 embeddings.
Evaluation Protocols. Our proposed TIPO leverages large-scale image caption datasets for training, which overlap with the training text distributions of many T2I models. Following Promptist [26], we divide our experiments into two settings: (1) In-domain, where the T2I model’s training texts overlap with those used by TIPO, and (2) Out-of-domain, where no overlap exists.
5.2 In-domain Tag-based Prompt Optimization
To assess prompt optimization performance on tag-based prompts, we generate scenery images, as they contain abundant descriptive tags for objects and backgrounds. We randomly sample 32,768 tag-based prompts from Danbooru2023 [69, 46], shuffle and concatenate the scenery tags into new prompts (thereby preventing data leakage of the original captions), and generate one image per prompt using Kohaku-XL-Zeta, an SDXL [50, 46] variant fine-tuned on Danbooru2023. This evaluation tests the in-domain capabilities, as Danbooru2023 is also used during TIPO training.
Metric Original GPT MagicPrompt Promptist TIPO (Ours)
FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
Vendi Score $\uparrow$ 16.8135 8.6632 11.9014 14.3273 13.3070
Table 1: Performance on in-domain tag-based prompts, with best results bolded and the second-best underlined.
The results in Table 1 reveal two key insights. First, MagicPrompt and Promptist, which rely on user prompts or reinforcement learning, underperform in Aesthetic and AI Corrupted Scores due to the quality or quantity limitations of their collected samples (e.g., Promptist uses 90K samples, limited by the high reinforcement learning cost). In contrast, GPT and TIPO benefit from large-scale training corpora ($>$30M samples), yielding higher-quality outputs. Second, TIPO achieves the best FDD by a substantial margin over GPT, which can be attributed to its superior distribution alignment with T2I models.
5.3 In-domain NL-based Prompt Optimization
We evaluate the prompt optimization performance on NL-based prompts by selecting 10,000 short prompts and 10,000 long prompts from CaptionEmporium [12] and GBC [29] as test prompts. In particular, since the long prompts are much longer than typical user input, we truncate them to two sentences ($<$ 40 words) to simulate real-world applications. We use SDXL-1.0-base [50] as the T2I model, whose training text data largely overlap with TIPO.
Short Original GPT MagicPrompt Promptist TIPO(Ours)
FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
Vendi Score $\uparrow$ 38.1715 34.7139 38.1553 34.1270 37.0649
Truncated Long Original GPT MagicPrompt Promptist TIPO(Ours)
FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
Vendi Score $\uparrow$ 38.2533 34.8108 37.8412 33.5266 37.0900
Table 2: Performance on in-domain NL-based prompts.
@@ 5.3 In-domain NL-based Prompt Optimization
Figure 5: Generated images from 4 types of prompts: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated ($<$ 40 words) long prompt, (d) TIPO-enhanced truncated prompt. TIPO adds detail and maintains variety, yielding coherent images from simple prompts.
Table 2 demonstrates that TIPO achieves either the best or second-best scores in Aesthetic and AI Corrupt Score by effectively enriching the original prompt with appropriate textual elements while rarely introducing extraneous noise. While all methods compromise fidelity and diversity, as reflected in the FDD and Vendi Score, TIPO remains competitive because it maintains small semantic deviation from the original sentences via progressive refinement.
(a) Oringial Prompt
(b) Text Completion
(c) TIPO
Figure 6: Qualitative comparison of prompt completion versus TIPO (ours). (a) shows images generated by original prompts. (b) depicts images generated by text completion using TIPO-500M. (c) shows images generated by TIPO.
5.4 Out-of-domain Performance
Some recent T2I models are trained on proprietary images and captions. As a representative example, SD-3.5-Large[20] is trained on private images captioned with CogView[71], which differ markedly from the texts used to train TIPO. To evaluate model performance in this out-of-domain scenario, we generate 8,192 original tag- and NL-based prompts using the baseline GPT-4o-mini rather than relying on existing prompt datasets. We apply the remaining methods to these prompts and assess their performance.
GPT MagicPrompt Promptist TIPO(Ours)
Aesthetic Score $\uparrow$ 6.7125 6.4507 6.3924 6.0536
AI Corrupt Score $\uparrow$ 0.9482 0.8577 0.9053 0.9280
Vendi Score $\uparrow$ 8.9718 15.8721 16.4891 21.5706
Table 3: Performance on the out-of-domain T2I model
As shown in Table 3, SD-3.5-Large faithfully generates images that align well with GPT-produced prompts. Consequently, additional optimizations tend to reduce fidelity and introduce more artifacts. Nevertheless, GPT-generated prompts are accurate and lack diversity. TIPO optimization enriches the prompts with additional details that harmonize with the original themes, significantly enhancing the diversity of the generated images.
5.5 Human Preference Evaluation
Quantitative metrics may not fully align with human preference. Therefore, we conducted a user study based on pairwise image comparisons between the original prompt, MagicPrompt, Promptist, and TIPO on over 1,400 images, gathering preferences from 221 volunteers. As illustrated in Figure 7, TIPO achieved the highest overall win rate at 51.3%, significantly outperforming competitors. In out-of-domain scenarios, TIPO’s win rate increased to 52.5%, demonstrating consistently strong user preference across different contexts. For further results and statistics, please refer to Appendix F.
Figure 7: The win-tie-lose rate on overall preference on all data and out-of-domain situation where images are generated by SD-3.5-Medium.
Figure 8: ELO rating on overall preference on all data
5.6 Ablation Study
Original Completion TIPO
Aesthetic Score $\uparrow$ 5.9507 5.9631 5.9960
Corrupt Score $\uparrow$ 0.6799 0.8131 0.8971
Vendi Score $\uparrow$ 19.7208 18.2896 13.9864
Table 4: Ablation results
We attribute TIPO’s success primarily to its multi-task pre-sampling design detailed in Section 4.3. However, a natural question arises: can a simple text completion approach achieve similar or better results? To validate our design, we compare TIPO’s pre-sampling strategy with a direct text completion using the same language model, TIPO-500M. As shown in Table 4, TIPO significantly improves aesthetic scores and reduces artifacts compared to text completion. For visual examples, please refer to Figure 6. For more ablation details and visualization, please refer to Appendix 5.6.
@@ 6 Conclusion
We introduced TIPO, a lightweight prompt pre-sampling framework designed for efficient real-world Text-to-Image (T2I) applications. By aligning user prompts with the intrinsic distributions of T2I training datasets, TIPO enhances semantic coherence, image fidelity, and diversity with minimal inference overhead.
TIPO consistently outperforms existing prompt optimization methods, achieving superior performance across multiple evaluation metrics. Its ability to preserve user intent while enhancing prompt diversity and image quality makes it well-suited for both in-domain and out-of-domain scenarios, demonstrating robustness across various prompt formats and model architectures.
Extensive user studies also confirmed TIPO’s strong alignment with human preferences, further highlighting its potential for practical applications. To encourage wider adoption and facilitate reproducibility, we release our trained models and source code. We hope TIPO will inspire further advancements in efficient, scalable, and robust generative frameworks for creative systems.